# InteractiveOmni: A Unified Omni-modal Model for Audio-Visual Multi-turn Dialogue

### v1

@@ 2.2.1 Image Data
GeneralQA $VQAv2$ , Visual7Wzhu2016visual7wgroundedquestionanswering , ViRL39KViRL39K_vl-rethinker , MMDUliu2024mmdumultiturnmultiimagedialog , VISTVIST_huang2016visual , GQAgqa , OKVQAmarino2019ok_okvqa
@@ 3.1.1 Multi-modal Multi-turn Memory Benchmark(MMMB)
We evaluate InteractiveOmni against a suite of state-of-the-art open-source and closed-source MLLMs using Gemini-2.5-Pro as the judge model for assessment comanici2025gemini . As shown in Table 6, InteractiveOmni-4B outperforms leading vision-language models such as Qwen2.5-VL-7B bai2025qwen25 and InternVL3-8B (zhu2025internvl3, ), as well as Omni-MLLMs including Qwen2.5-Omni-7B (xu2025qwen2, ) and GPT‑4o-mini (gpt4o, ). InteractiveOmni-8B further strengthens the multi-turn performance and is comparable to Gemini-2.5-Flash  (comanici2025gemini, ) (58.17 vs. 60.84), demonstrating the strong performance in long-term memory for historical image and text context. To quantify the model performance degradation related to memory information, we conduct two types of evaluation: one measures accuracy based on the number of historical images to be recalled, and the other assesses accuracy according to the turn distance between critical historical turns and the final question. As shown in Figure 5, the model’s performance decreases as the turn distance increases. InteractiveOmni-4B maintains an accuracy of 40% even with a turn distance of four. This demonstrates its robustness, which is comparable to the proprietary model such as Gemini-2.5-Flash, and significantly surpasses other open-source models like Qwen2.5-Omni-7B, InternVL3-8B, and Qwen2.5-VL-7B, which only achieve a score of around 20. Additionally, all models exhibit a significant performance decline as the number of images to be memorized increases, highlighting a common weakness in the long-term memory of current MLLMs. For example, even Gemini-2.5-Flash achieves a score of only 20 under these conditions.
@@
Proprietary GPT‑4o-mini (gpt4o, ) 70.00 29.41 58.06 51.33
Gemini-2.5-Flash (comanici2025gemini, ) 75.76 40.19 70.97 60.84
Open-source InternVL3-8B (zhu2025internvl3, ) 31.31 15.69 33.87 25.86
Qwen2.5-VL-7B (bai2025qwen25, ) 35.35 13.73 27.42 25.10
Qwen2.5-Omni-7B (xu2025qwen2, ) 32.32 9.80 40.32 25.48
@@ 3.2.4 Omni-modal Understanding Benchmarks
We evaluate the omni-modal benchmark OmniBench (li2024omnibench, ) to assess the omni-modal understanding capability of MLLMs, and compare InteractiveOmni with Qwen2.5-Omni-7B and other models. As shown in Table 12, InteractiveOmni-4B achieves state-of-the-art performance on OmniBench, attaining an average score of 59.19 that substantially exceeds other Omni models, demonstrating exceptional omni-modal understanding capabilities.
@@
Gemini-1.5-Pro (team2024gemini, ) 42.67 42.26 46.23 42.91
MIO-Instruct (wang2024mio, ) (7B) 36.96 33.58 11.32 33.80
AnyGPT (7B) (zhan2024anygpt, ) 17.77 20.75 13.21 18.04
video-SALMONN (13B) (sun2024video, ) 34.11 31.70 56.60 35.64
UnifiedIO2-xlarge (3.2B) (lu2024unified, ) 39.56 36.98 29.25 38.00
UnifiedIO2-xxlarge (6.8B) (lu2024unified, ) 34.24 36.98 24.53 33.98
MiniCPM-o-2.6 (minicpm-v, ) - - - 40.50
Baichuan-Omni-1.5 (li2025baichuan, ) - - - 42.90
Qwen2.5-Omni-7B (xu2025qwen2, ) 55.25 60.00 52.83 56.13
@@
Qwen2-Audio (chu2024qwen2, ) 42.77 | 69.67 | 45.20 | 40.30 | 57.19 | 51.03
GLM-4-Voice (zeng2024glm4voiceintelligenthumanlikeendtoend, ) 47.43 | 76.00 | 55.40 | 51.80 | 57.89 | 57.70
VITA-1.5 (fu2025vita, ) 41.00 | 74.20 | 57.30 | 46.80 | 68.20 | 57.50
Step-Audio-chat (huang2025stepaudiounifiedunderstandinggeneration, ) 60.00 | 72.33 | 73.00 | 56.80 | 56.53 | 63.73
Baichuan-Audio (li2025baichuanaudiounifiedframeworkendtoend, ) 41.90 | 78.40 | 64.50 | 61.70 | 77.40 | 64.78
Kimi-Audio (kimiteam2025kimiaudiotechnicalreport, ) 58.02 | 79.33 | 70.20 | 62.10 | 75.73 | 69.08
MiniCPM-o-2.6 (minicpm-v, ) 38.60 | 77.80 | 68.60 | 61.90 | 51.80 | 59.74
Baichuan-Omni-1.5 (li2025baichuan, ) 50.00 | 78.50 | 59.10 | 57.20 | 77.90 | 64.54
Qwen2.5-Omni-7B (xu2025qwen2, ) 63.76 | 75.33 | 62.80 | 57.06 | 72.76 | 66.34
@@
Qwen2-Audio (chu2024qwen2, ) 3.69 | 3.40 | 3.01 | 35.35 | 35.43
GLM-4-Voice (zeng2024glm4voiceintelligenthumanlikeendtoend, ) 4.06 | 3.48 | 3.18 | 43.31 | 40.11
VITA-1.5 (fu2025vita, ) 4.21 | 3.66 | 3.48 | 38.88 | 52.15
Step-Audio-chat (huang2025stepaudiounifiedunderstandinggeneration, ) 3.99 | 2.99 | 2.93 | 46.84 | 28.72
Baichuan-Audio (li2025baichuanaudiounifiedframeworkendtoend, ) 4.41 | 4.08 | 3.92 | 45.84 | 53.19
Kimi-Audio (kimiteam2025kimiaudiotechnicalreport, ) 4.46 | 3.97 | 4.20 | 63.12 | 62.17
MiniCPM-o-2.6 (minicpm-v, ) 4.42 | 4.15 | 3.94 | 50.72 | 54.78
Baichuan-Omni-1.5 (li2025baichuan, ) 4.50 | 4.05 | 4.06 | 43.40 | 57.25
Qwen2.5-Omni-7B (xu2025qwen2, ) 4.50 | 3.84 | 3.89 | 56.40 | 61.32
@@
Qwen2-Audio (chu2024qwen2, ) 49.01 | 54.70 | 22.57 | 98.85 | 55.32
GLM-4-Voice (zeng2024glm4voiceintelligenthumanlikeendtoend, ) 52.97 | 52.80 | 24.91 | 88.08 | 57.40
VITA-1.5 (fu2025vita, ) 71.65 | 55.30 | 38.14 | 97.69 | 64.53
Step-Audio-chat (huang2025stepaudiounifiedunderstandinggeneration, ) 31.87 | 50.60 | 29.19 | 65.77 | 50.13
Baichuan-Audio (li2025baichuanaudiounifiedframeworkendtoend, ) 71.65 | 54.80 | 50.31 | 99.42 | 69.27
Kimi-Audio (kimiteam2025kimiaudiotechnicalreport, ) 83.52 | 69.70 | 61.10 | 100.0 | 76.91
MiniCPM-o-2.6 (minicpm-v, ) 78.02 | 60.40 | 49.25 | 97.69 | 71.23
Baichuan-Omni-1.5 (li2025baichuan, ) 74.51 | 62.70 | 54.54 | 97.31 | 71.32
Qwen2.5-Omni-7B (xu2025qwen2, ) 80.90 | 66.70 | 53.50 | 99.20 | 73.60
@@ 3.2.5 Speech-to-text Benchmarks
To assess the speech understanding and speech-based question-answering capabilities, we evaluate InteractiveOmni on the following benchmarks: OpenAudioBench (li2025baichuanaudiounifiedframeworkendtoend, ) and VoiceBench (chen2024voicebench, ). As shown in Table 13, our model performs excellently on the speech-to-text question-answering benchmarks, outperforming recent open-source audio language models and omni models. InteractiveOmni-4B achieves an average score of 69.10 on OpenAudioBench, significantly outperforming Kimi-Audio kimiteam2025kimiaudiotechnicalreport , Step-Audio-chat huang2025stepaudiounifiedunderstandinggeneration and Qwen2.5-Omni-7B xu2025qwen2 . These voice-chat benchmarks reflect our model’s substantial progress in diversified speech interaction.

### v2

@@ 2.2.1 Image Data
GeneralQA VQAv2VQAv2, Visual7Wzhu2016visual7wgroundedquestionanswering, ViRL39KViRL39K_vl-rethinker, MMDUliu2024mmdumultiturnmultiimagedialog, VISTVIST_huang2016visual, GQAgqa, OKVQAmarino2019ok_okvqa
@@ 3.1.1 Multi-modal Multi-turn Memory Benchmark(MMMB)
We evaluate InteractiveOmni against a suite of state-of-the-art open-source and closed-source MLLMs using Gemini-2.5-Pro as the judge model for assessment comanici2025gemini. As shown in Table 6, InteractiveOmni-4B outperforms leading vision-language models such as Qwen2.5-VL-7B bai2025qwen25 and InternVL3-8B (zhu2025internvl3), as well as Omni-MLLMs including Qwen2.5-Omni-7B (xu2025qwen2) and GPT‑4o-mini (gpt4o). InteractiveOmni-8B further strengthens the multi-turn performance and is comparable to Gemini-2.5-Flash  (comanici2025gemini) (58.17 vs. 60.84), demonstrating the strong performance in long-term memory for historical image and text context. To quantify the model performance degradation related to memory information, we conduct two types of evaluation: one measures accuracy based on the number of historical images to be recalled, and the other assesses accuracy according to the turn distance between critical historical turns and the final question. As shown in Figure 5, the model’s performance decreases as the turn distance increases. InteractiveOmni-4B maintains an accuracy of 40% even with a turn distance of four. This demonstrates its robustness, which is comparable to the proprietary model such as Gemini-2.5-Flash, and significantly surpasses other open-source models like Qwen2.5-Omni-7B, InternVL3-8B, and Qwen2.5-VL-7B, which only achieve a score of around 20. Additionally, all models exhibit a significant performance decline as the number of images to be memorized increases, highlighting a common weakness in the long-term memory of current MLLMs. For example, even Gemini-2.5-Flash achieves a score of only 20 under these conditions.
@@
Proprietary GPT‑4o-mini (gpt4o) 70.00 29.41 58.06 51.33
Gemini-2.5-Flash (comanici2025gemini) 75.76 40.19 70.97 60.84
Open-source InternVL3-8B (zhu2025internvl3) 31.31 15.69 33.87 25.86
Qwen2.5-VL-7B (bai2025qwen25) 35.35 13.73 27.42 25.10
Qwen2.5-Omni-7B (xu2025qwen2) 32.32 9.80 40.32 25.48
@@ 3.2.4 Omni-modal Understanding Benchmarks
We evaluate the omni-modal benchmark OmniBench (li2024omnibench) to assess the omni-modal understanding capability of MLLMs, and compare InteractiveOmni with Qwen2.5-Omni-7B and other models. As shown in Table 12, InteractiveOmni-4B achieves state-of-the-art performance on OmniBench, attaining an average score of 59.19 that substantially exceeds other Omni models, demonstrating exceptional omni-modal understanding capabilities.
@@
Gemini-1.5-Pro (team2024gemini) 42.67 42.26 46.23 42.91
MIO-Instruct (wang2024mio) (7B) 36.96 33.58 11.32 33.80
AnyGPT (7B) (zhan2024anygpt) 17.77 20.75 13.21 18.04
video-SALMONN (13B) (sun2024video) 34.11 31.70 56.60 35.64
UnifiedIO2-xlarge (3.2B) (lu2024unified) 39.56 36.98 29.25 38.00
UnifiedIO2-xxlarge (6.8B) (lu2024unified) 34.24 36.98 24.53 33.98
MiniCPM-o-2.6 (minicpm-v) - - - 40.50
Baichuan-Omni-1.5 (li2025baichuan) - - - 42.90
Qwen2.5-Omni-7B (xu2025qwen2) 55.25 60.00 52.83 56.13
@@
Qwen2-Audio (chu2024qwen2) 42.77 | 69.67 | 45.20 | 40.30 | 57.19 | 51.03
GLM-4-Voice (zeng2024glm4voiceintelligenthumanlikeendtoend) 47.43 | 76.00 | 55.40 | 51.80 | 57.89 | 57.70
VITA-1.5 (fu2025vita) 41.00 | 74.20 | 57.30 | 46.80 | 68.20 | 57.50
Step-Audio-chat (huang2025stepaudiounifiedunderstandinggeneration) 60.00 | 72.33 | 73.00 | 56.80 | 56.53 | 63.73
Baichuan-Audio (li2025baichuanaudiounifiedframeworkendtoend) 41.90 | 78.40 | 64.50 | 61.70 | 77.40 | 64.78
Kimi-Audio (kimiteam2025kimiaudiotechnicalreport) 58.02 | 79.33 | 70.20 | 62.10 | 75.73 | 69.08
MiniCPM-o-2.6 (minicpm-v) 38.60 | 77.80 | 68.60 | 61.90 | 51.80 | 59.74
Baichuan-Omni-1.5 (li2025baichuan) 50.00 | 78.50 | 59.10 | 57.20 | 77.90 | 64.54
Qwen2.5-Omni-7B (xu2025qwen2) 63.76 | 75.33 | 62.80 | 57.06 | 72.76 | 66.34
@@
Qwen2-Audio (chu2024qwen2) 3.69 | 3.40 | 3.01 | 35.35 | 35.43
GLM-4-Voice (zeng2024glm4voiceintelligenthumanlikeendtoend) 4.06 | 3.48 | 3.18 | 43.31 | 40.11
VITA-1.5 (fu2025vita) 4.21 | 3.66 | 3.48 | 38.88 | 52.15
Step-Audio-chat (huang2025stepaudiounifiedunderstandinggeneration) 3.99 | 2.99 | 2.93 | 46.84 | 28.72
Baichuan-Audio (li2025baichuanaudiounifiedframeworkendtoend) 4.41 | 4.08 | 3.92 | 45.84 | 53.19
Kimi-Audio (kimiteam2025kimiaudiotechnicalreport) 4.46 | 3.97 | 4.20 | 63.12 | 62.17
MiniCPM-o-2.6 (minicpm-v) 4.42 | 4.15 | 3.94 | 50.72 | 54.78
Baichuan-Omni-1.5 (li2025baichuan) 4.50 | 4.05 | 4.06 | 43.40 | 57.25
Qwen2.5-Omni-7B (xu2025qwen2) 4.50 | 3.84 | 3.89 | 56.40 | 61.32
@@
Qwen2-Audio (chu2024qwen2) 49.01 | 54.70 | 22.57 | 98.85 | 55.32
GLM-4-Voice (zeng2024glm4voiceintelligenthumanlikeendtoend) 52.97 | 52.80 | 24.91 | 88.08 | 57.40
VITA-1.5 (fu2025vita) 71.65 | 55.30 | 38.14 | 97.69 | 64.53
Step-Audio-chat (huang2025stepaudiounifiedunderstandinggeneration) 31.87 | 50.60 | 29.19 | 65.77 | 50.13
Baichuan-Audio (li2025baichuanaudiounifiedframeworkendtoend) 71.65 | 54.80 | 50.31 | 99.42 | 69.27
Kimi-Audio (kimiteam2025kimiaudiotechnicalreport) 83.52 | 69.70 | 61.10 | 100.0 | 76.91
MiniCPM-o-2.6 (minicpm-v) 78.02 | 60.40 | 49.25 | 97.69 | 71.23
Baichuan-Omni-1.5 (li2025baichuan) 74.51 | 62.70 | 54.54 | 97.31 | 71.32
Qwen2.5-Omni-7B (xu2025qwen2) 80.90 | 66.70 | 53.50 | 99.20 | 73.60
@@ 3.2.5 Speech-to-text Benchmarks
To assess the speech understanding and speech-based question-answering capabilities, we evaluate InteractiveOmni on the following benchmarks: OpenAudioBench (li2025baichuanaudiounifiedframeworkendtoend) and VoiceBench (chen2024voicebench). As shown in Table 13, our model performs excellently on the speech-to-text question-answering benchmarks, outperforming recent open-source audio language models and omni models. InteractiveOmni-4B achieves an average score of 69.10 on OpenAudioBench, significantly outperforming Kimi-Audio kimiteam2025kimiaudiotechnicalreport, Step-Audio-chat huang2025stepaudiounifiedunderstandinggeneration and Qwen2.5-Omni-7B xu2025qwen2. These voice-chat benchmarks reflect our model’s substantial progress in diversified speech interaction.